OpenStudy (notsmartatall):

determine the equation whose roots are -3, -3 and 3

OpenStudy (mathstudent55):

Any root \(a\) of a polynomial equation comes form a factor \(x - a\).

OpenStudy (notsmartatall):

a. x^3+3x^2-9x-27=0 b. x^3-9x^2+27x+8=0 c. x^3-3x^2+9x=0 d. x^3-3x^2+8x=0

OpenStudy (anonymous):

Well which one can you eliminate right away?

OpenStudy (mathstudent55):

\( [x - (-3)][x - (-3)][x - 3] \) \( = (x + 3)(x + 3)(x - 3) \) Look at the last terms: 3, 3, -3. When you multiply then together, what do you get?

OpenStudy (anonymous):

Hint: you can get rid of the answers with even numbers in them. :D

OpenStudy (notsmartatall):

so its c

OpenStudy (anonymous):

Work it out.

OpenStudy (mathstudent55):

Did you find the product of 3, 3, and -3?

OpenStudy (anonymous):

(x-3)(x+3)^2 what gives you this expression?

OpenStudy (notsmartatall):

ahhh @iambatman your not helping!! lol jk how do you set it up first

OpenStudy (anonymous):